OpenStudy (anonymous):

A regulation tennis court for a doubles match is laid out so that its length is 6 feet more than two times its width. The area of the doubles court is 2808 square feet. Find the length and width of a doubles court.

OpenStudy (mathstudent55):

let width = w length = 2w + 6 area = LW = 2808 (2w + 6)w = 2808 Can you continue now?

OpenStudy (anonymous):

So...2w^+6w = 2808

OpenStudy (anonymous):

But will I end up having to use the quad. formula?

OpenStudy (anonymous):

B/c I've no idea how to solve this using factoring.

OpenStudy (mathstudent55):

first term is 2w^2. yes you'll need the quadratic formula or factoring

OpenStudy (anonymous):

Sorry, that was a typo.

OpenStudy (mathstudent55):

I thought so. First factor out a 2

OpenStudy (anonymous):

w^2+3 = 1404

OpenStudy (anonymous):

Argh...3w

OpenStudy (mathstudent55):

w^2 + 3w - 1404 = 0

OpenStudy (anonymous):

Right. So x = -3 +/- sqrt(9 - 4(1)(-1404))/2

OpenStudy (anonymous):

x = -3 +/- sqrt(9 - 5616)/2

OpenStudy (anonymous):

Okay...there's no way that's right. What have I done? lol

OpenStudy (mathstudent55):

In the root, you have -4(1)(-1404) which is positive, not negative

OpenStudy (anonymous):

Okay, then...x = -3 +/- sqrt(5626)/2

OpenStudy (mathstudent55):

right

OpenStudy (anonymous):

x = (-3 +/- 75)/2 -> -39 or 36

OpenStudy (anonymous):

So w = 36 and l = 78

OpenStudy (mathstudent55):

Good. Now -39 is meaningless as a width of a rectangle, so discard it. You have w = 36 L = 2w + 6 =...

OpenStudy (mathstudent55):

right

OpenStudy (anonymous):

Okay...thank you once again!!

OpenStudy (mathstudent55):

wlcm